Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
cifaldi, nardecchia - 18:01 Friday 17 November 2023 (62485) Print this report
TCS configuration for the weekend shifts

According to the commissioning crew, after the daily meeting, we settled the PR CHRoCC, the NI DAS IN and OUT powers in preparation for the weekend activity (PR CHRoCC and SR RH scans with the ITF at CITF configuration). The starting value of the PR CHRoCC has been set to 0.4 V at 16.20 UTC with a ramp of 1.5 hr. 

At 15.27 UTC NI DAS IN and OUT needed to compensate for the cold lens (INNER 0.08 W, OUTER 1.8 W) have been settled.

The CH flip mirrors (REL3 and REL7) have been closed and Piernicola commented on the lines in the automation file to avoid to re-open them when the ITF unlocks.

Comments to this report:
nardecchia - 15:09 Sunday 19 November 2023 (62494) Print this report

This morning we realized that we forgot to re-open the flip mirror of the NI Inner (REL 5 ) after settled the right power on Friday afternoon.

The flip mirror has been re-opened today, at 11.31 UTC.


Search Help